Find the first five terms of the sequence defined by each of these recurrence relations and initial conditions.

Then solve the recurrence relation. a) an = an-1 + 3, a 0 = 3
Answer #1

First Five Terms

a0 = 3

a1 = a0 + 3 = 3 + 3 = 6

a2 = a1 + 3 = 6 + 3 = 9

a3 = a2 + 3 = 9 + 3 = 12

a4 = a3 + 3 = 12 + 3 = 15

Formula

an = 3*(n+1)

Explanation

We can observe that these all are multiples of 3

Please up vote